OpenStudy (anonymous):

Show the conjecture is false by finding a counterexample. The product of (a+b)^2 is equal to a^2 +b^2, for a is not equal to 0, and b is not equal to 0.

OpenStudy (anonymous):

Say, a=3 and b=4 (3+4)^2=49 3^2+4^2=9+16=25 Which is not equal to 49.